# Explain Like I’m Five: Using LLMs to Improve PDE Surrogate Models with Text

### v4

@@ 5.4 Ablation Study
Lastly, we perform an ablation study based on text information, to determine which components of our text description are most helpful during the learning process. We compare our baseline model against our multimodal model with SentenceTransformer using 4,000 samples for each equation. Both baseline and multimodal models were trained with transfer learning. Results for next-step prediction are given in figure 5, and autoregressive accumulated rollout errors are given in 2. In both experiments, we see a correlation between richness of sentence description and performance, but that every level of text information improves upon baseline model with transfer learning. Additional error plots and results are given in appendix C.
| Transfer | None | B | C | Q | BC | BQ | CQ | BCQ | |
| Combined | 1.30 0.08 | 0.84 0.02 | 0.85 0.01 | 0.85 0.08 | 0.87 0.01 | 0.82 0.07 | 0.80 0.02 | 0.82 0.07 | 0.80 0.09 |
| Burgers | 4.06 3.93 | 1.02 0.04 | 0.86 0.08 | 1.10 0.05 | 1.04 0.07 | 0.86 0.04 | 0.86 0.09 | 1.00 0.07 | 0.85 0.07 |
| Heat | 1.54 0.27 | 0.37 0.02 | 0.30 0.19 | 0.38 0.03 | 0.38 0.05 | 0.20 0.04 | 0.19 0.03 | 0.38 0.02 | 0.30 0.17 |
| Navier Stokes | 0.85 0.57 | 0.19 0.15 | 0.29 0.04 | 0.26 0.02 | 0.33 0.06 | 0.27 0.02 | 0.27 0.01 | 0.24 0.03 | 0.25 0.00 |
| Shallow Water | 52.20 50.63 | 82.06 33.18 | 10.02 9.98 | 56.15 50.17 | 51.89 33.97 | 63.49 50.75 | 43.81 50.04 | 65.07 75.05 | 51.22 60.16 |
@@ 6 Discussion
The aim of our multimodal approach is to incorporate known system information through sentence descriptions, rather than complex data conditioning strategies. We see for our next-step prediction and autoregressive rollout tasks that a multimodal approach is able to significantly improve performance over baseline. SentenceTransformer tends to improve performance more than Llama, and this is likely due to the nature of the embeddings generated by each LLM. SentenceTransformer is trained to generate useful embeddings at the sentence level, whereas Llama is trained to generate word-level embeddings. In order to utilize Llama for out framework, we need to average over all word-level embeddings, which can lose subtle information in our embeddings. The sentence-level embeddings from SentenceTransformer, on the other hand, allows us to use these embeddings directly from the LLM in our multimodal cross-attention block.
We can visualize the embeddings from SentenceTransformer for each level of text description as well in our combined dataset. In figure 6, we have t-SNE embeddings for sentences that simply identify the equation compared against t-SNE embeddings with our full text description, that is with boundary condition, operator coefficient, and qualitative information. In our combined data set we only have samples from the Heat, Burgers, and Navier-Stokes equations, which is reflected in the three clusters seen. However, with our full text description, we have a significantly richer embedding. We see large clusters with subregions with our Heat and Burgers samples, matching our continuous distributions for operator coefficients and boundary condition values [analyze further?]. This shows how we are able to easily introduce significantly more structure into our data, making the learning task easier. t-SNE plots were generated with Chan et al. (2019).
@@ 7 Conclusion
We introduced a multimodal approach to PDE surrogate modeling that utilizes LLMs and sentence descriptions of our systems that significantly improves performance for a variety of tasks. Our multimodal approach allows us to easily incorporate system information that both captures quantitative and qualitative aspects of our governing equations. Analysis of our sentence embeddings shows that we are able to capture increasing amounts of underlying structure in our data by simply adding more information through text, rather than introducing additional mathematical approaches, such as constrained loss functions that are popular with PINNs.
While this direction is promising, performance could be further improved with tuning out LLMs directly, rather than just training an embedding layer that uses LLM output, and we leave this to future work. Additionally, pretraining strategies have shown success in various LLM applications from natural language processing to PDE surrogate modeling. While this does introduce significant computational overhead, it may improve performance and is also a potential future direction to take. Lastly, we only benchmarked the FactFormer, but this framework and multimodal block is not specific to FactFormer, and may prove useful for improving other physics-based models such as FNO and DeepONet.
@@ Appendix A Sentence Descriptions
| Burgers: | Burgers equation models a conservative system that can develop shock wave discontinuities. Burgers equation is a first order quasilinear hyperbolic partial differential equation. |
@@
Appendix B Hyperparameters
B.1 Architecture
We use identical architecture across all of our experiments, given in table 3. We note that the hidden dimension controls both the FactFormer embedding dimension as well as the LLM projection head embedding dimension. This is done so we can use our multimodal cross-attention block without any additional data projections.
@@ B.1 Architecture
| FactFormer | 1 | 128 | 64 | 4 | 2 | 2 |
Our convolutional patch embedding used a kernel size of 8, and our convolutional upsampling layer used a ConvTranspose layer with kernel size of 8 with a stride size of 4, followed by two fully connected layers with widths of 128, and GELU activation. Additionally, our sentence embedding projection layer used four fully-connected layers with size [384/4096, 256, 256, 128, no. patches] to match the number of patches from our patch embedding, with ReLU activation, then upsamples our channel dimension of 1 to match our hidden dimension of 128 using two fully connected layers with width 128 and GELU activation.
B.2 Training
Training hyperparameters were generally kept as similar as possible, although instability in training was noted for our Multimodal FactFormer at times during finetuning. Generally, lower learning rate and higher weight decay led to more stable training. Table 4 has training hyperparameters for our 4000 samples/equation case. We trained each model for 1000 epochs for each experiment. In each epoch, each trajectory is used, where the specific input/output frame for each trajectory is randomly sampled.
| Model | Batch Size | Learning Rate | Weight Decay | Finetune Learning Rate | Finetune Weight Decay |
| FactFormer | 64 | 1e-4 | 1e-8 | 1e-4 | 1e-8 |
| FactFormer (Transfer) | 64 | 1e-4 | 1e-8 | 1e-4 | 1e-8 |
| FactFormer + ST | 64 | 1e-4 | 1e-8 | 1e-4 | 1e-8 |
| FactFormer + ST (Transfer) | 64 | 1e-4 | 1e-8 | 1e-4 | 1e-8 |
| FactFormer + Llama | 64 | 5e-4 | 1e-8 | 5e-5 | 1e-8 |
| FactFormer + Llama (Transfer) | 64 | 5e-4 | 1e-8 | 5e-5 | 1e-8 |
Additional tuning was necessary for training stability for FactFormer + Llama with small amounts of data. Hyperparameters are given in table 5. No additional tuning was necessary for baseline FactFormer of FactFormer + ST.
| Samples/Equation | Batch Size | Learning Rate | Weight Decay | Finetune Learning Rate | Finetune Weight Decay |
| 50 | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 100 | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 500 | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 1000 | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 2000 | 256 | 5e-4 | 1e-8 | 5e-5 | 1e-8 |
| 4000 | 64 | 5e-4 | 1e-8 | 5e-5 | 1e-8 |
| 50 (Transfer) | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 100 (Transfer) | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 500 (Transfer) | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 1000 (Transfer) | 128 | 5e-4 | 1e-7 | 1e-4 | 1e-7 |
| 2000 (Transfer) | 256 | 5e-4 | 1e-8 | 5e-5 | 1e-8 |
| 4000 (Transfer) | 64 | 5e-4 | 1e-8 | 5e-5 | 1e-8 |
Appendix C Additional Next-Step and Rollout Results
| Baseline | ST | Llama | Transfer | ST + Transfer | Llama + Transfer | |
| Combined | 2.95 0.04 | 2.27 0.02 | 2.52 0.01 | 2.94 0.09 | 2.36 0.09 | 2.52 0.02 |
| Burgers | 2.93 0.02 | 1.95 0.08 | 2.53 0.13 | 2.63 0.13 | 1.81 0.03 | 1.96 0.05 |
| Heat | 1.23 0.04 | 0.73 0.04 | 0.95 0.03 | 1.12 0.03 | 0.81 0.05 | 0.55 0.11 |
| Navier Stokes | 1.57 0.04 | 1.75 0.16 | 2.45 0.01 | 1.80 0.21 | 1.97 0.02 | 2.08 0.02 |
| Shallow Water | 0.18 0.03 | 0.14 0.02 | 0.33 0.06 | 0.23 0.02 | 0.15 0.01 | 0.39 0.08 |
@@ Appendix C Additional Next-Step and Rollout Results
Figure 8 plots the autoregressive rollout error against timestep. Table 1 presents the sum over timesteps of these plots. Errors are first averaged over each sample for a given random seed, then the mean and standard deviation across seeds is plotted.
Appendix D Ablation Study Additional Results
| Baseline | None | B | C | Q | BC | BQ | CQ | BCQ | |
| Combined | 3.22 0.05 | 2.86 0.02 | 2.93 0.13 | 2.81 0.07 | 2.89 0.09 | 2.86 0.06 | 2.88 0.06 | 2.85 0.01 | 2.86 0.04 |
| Burgers | 3.13 0.17 | 2.35 0.11 | 2.32 0.20 | 2.32 0.12 | 2.44 0.16 | 2.31 0.21 | 2.35 0.10 | 2.33 0.12 | 2.34 0.12 |
| Heat | 1.30 0.07 | 1.00 0.06 | 0.97 0.06 | 1.00 0.07 | 1.02 0.08 | 0.95 0.06 | 0.96 0.05 | 1.00 0.06 | 0.96 0.06 |
| Navier Stokes | 2.01 0.07 | 2.20 0.02 | 2.20 0.01 | 2.19 0.03 | 2.23 0.14 | 2.20 0.02 | 2.12 0.02 | 2.11 0.03 | 2.12 0.03 |
| Shallow Water | 0.23 0.01 | 0.16 0.00 | 0.16 0.00 | 0.17 0.01 | 0.17 0.02 | 0.15 0.01 | 0.16 0.01 | 0.16 0.01 | 0.16 0.01 |
Figure 10 plots the autoregressive rollout error against timestep. Table 2 presents the sum over timesteps of these plots. Errors are first averaged over each sample for a given random seed, then the mean and standard deviation across seeds is plotted.
Appendix E Data Scaling Additional Results
Data scaling results for the combined data sets are given in figure 11, for the Navier-Stokes data set is given in 12, and for the Shallow Water equations is given in 13.
Appendix F t-SNE Additional Results
@@ Appendix E Data Scaling Additional Results
We see as we increase sentence complexity, we get additional structure in our t-SNE embeddings. Adding coefficient information captures the distribution of coefficients well, seen in both figure 14 and figure 15.
Appendix G Corrected Burgers’ Equation Description
Using an equation description of Burgers’ equation that accurately describes the viscid case as:
we run a the experimental setup from section 5.1 using 100 samples from each equation and otherwise identical hyperparameters. Results are given in figures 16 and 17. We find no significant difference between system descriptions, likely due to the fact that this description is constant for every sample of Burgers’ equation, and therefore does not provide much useful information.

### v5

@@ 5.4 Ablation Study
Lastly, we perform an ablation study based on text information to determine which components of our text description are most helpful during the learning process using our SentenceTransformer variant. Results are given in figure 6. An additional ablation study with our Llama variant is given in appendix C.4. The same experimental setup is used as in section 5.3. We see that performance correlates strongly with relevant sentence information. Only the viscosity and forcing term amplitude vary in our Navier Stokes data set, with periodic boundary conditions in all samples. We see that adding boundary condition information has no impact on performance. Similarly, in the Shallow Water data set, we have the same boundary conditions and flow conditions for each sample, and so there is a very weak correlation with additional system information. There is also a weak performance improvement with the addition of qualitative information. Qualitative information acts as an imperfect proxy for coefficient information for the Heat, Burgers, and Navier Stokes data sets. This leads to a small improvement in performance when it is present without coefficient information. However, when coefficient information is provided, we generally do not see much additional improvement in performance. With our Heat data set, we see improvement as more information is added, with full system information having best performance. However, with Burgers equation, we see weaker correlation, suggesting that the LLM is unable to learn more complex systems and distribution of values as well as the simpler heat equation, or discrete Navier Stokes coefficients.
@@ 6 Discussion
The aim of our multimodal approach is to incorporate known system information through sentence descriptions, rather than complex data conditioning strategies. We see for our next-step prediction and autoregressive rollout tasks the SentenceTransformer variant tends to perform best, but on the fixed-future task Llama performs best, while the tokenizer variant is between the two in performance. Although Llama is larger than SentenceTransformer, these results do not show a clear trend that Llama’s capacity is being utilized for this task. Additionally, our tokenizer variant having comparable performance to both pretrained LLMs suggests the semantic knowledge of these LLMs does not affect performance. This is also seen in the ablation study results, where qualitative information has a weak effect on performance, and almost no effect when coefficient information is present.
We can visualize the embeddings from SentenceTransformer and Llama for each level of text description as well in our combined dataset. In figure 7, we have t-SNE embeddings for sentences with the full text description, that is with boundary condition, operator coefficient, and qualitative information. t-SNE plots for all other combinations of information are given in appendix E. For both SentenceTransformer and Llama, we have clusters corresponding to combinations of boundary conditions and governing equation This is despite that we average over sequence length to obtain Llama text embeddings. While we see clear clusters for equations and boundary conditions, we do not see mixed clusters seen in PICL (Lorsung & Farimani, 2024). These clusters do not match intuition that diffusion dominated Burgers equation behaves similarly to the Heat equation, suggesting that these pretrained LLMs lack semantic understanding of PDE systems. t-SNE plots were generated with TSNE-CUDA (Chan et al., 2019).
@@ 7 Conclusion
We introduced a multimodal approach to PDE surrogate modeling that utilizes LLMs and sentence descriptions of our systems that significantly improves performance for a variety of tasks. Our multimodal approach allows us to easily incorporate system information that both captures quantitative and qualitative aspects of our governing equations. Analysis of our sentence embeddings shows that we are able to capture increasing amounts of underlying structure in our data by simply adding more information through text, rather than introducing additional mathematical approaches, such as constrained loss functions that are popular with PINNs. However, marginal improvement when using qualitative information, and comparable performance to a simple tokenizer suggests that the semantic understanding of LLMs is not being fully utilized in this case.
While this direction is promising, performance could be further improved with tuning LLMs directly, rather than just training an embedding layer that uses LLM output. PDEs are defined rigorously and precisely, and so training LLMs specifically for PDE applications may also allow the semantic understanding of LLMs to be used in similar settings. We leave this to future work. Additionally, pretraining strategies have shown success in various LLM applications from natural language processing to PDE surrogate modeling. While this does introduce significant computational overhead, it may improve performance and is also a potential future direction to take. Lastly, we only benchmarked the FactFormer, but this framework and multimodal block is not specific to FactFormer, and may prove useful for improving other popular PDE surrogate models such as FNO and DeepONet.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Sentence Descriptions
Burgers: | Burgers equation models a conservative system that can develop shock wave discontinuities. Burgers equation is a second order partial differential equation. |
@@
Appendix B Training Setup
B.1 Model Architecture
Model architecture was kept the same for the multimodal FactFormer variants for each experiment. We note that the hidden dimension controls both the FactFormer embedding dimension as well as the LLM projection head embedding dimension. This is done so we can use our multimodal cross-attention block without any additional data projections. Our convolutional patch embedding used a kernel size of 8, and our convolutional upsampling layer used a ConvTranspose layer with kernel size of 8 with a stride size of 4, followed by two fully connected layers with widths of 128, and GELU activation. Additionally, our sentence embedding projection layer used a single hidden layer of size [384/4096/4096, no. patches] project the embedding dimension of SentenceTransformer, Llama, and Llema, respectively, down to match the number of patches from our patch embedding. The embeddings are then upsamples our channel dimension of 1 to match our hidden dimension using a single layer.
B.2 Next-Step
@@
FactFormer | 1 | 32 | 4 | 4 | 2 | 2 |
Multimodal FactFormer | 1 | 32 | 1 | 1 | 2 | 2 |
B.3 Fixed-Future
FactFormer | 1 | 64 | 4 | 4 | 2 | 2 |
Multimodal FactFormer | 1 | 32 | 4 | 4 | 2 | 2 |
B.4 Hyperparameters
Training hyperparameters such as batch size, learning rate, and weight decay were hand tuned for each experiment to reduce error. All models were trained with the Adam optimizer (Kingma & Ba, 2017).
Appendix C Additional Training Results
Results for our combined data set, numerical values for each experiment, and the ablation study with Llama as the LLM backbone are presented here. For all results, bold indicates the lowest error.
C.1 Additional Next-Step Results
Table 3 below gives numerical values for figure 8. These are the same results as in figure 3 with results for our combine data set included. We see that the combined data set has a similar trend to our Navier Stokes data set.
Baseline | ST | Tokenizer | Llama | Transfer | ST + Transfer | Tokenizer + Transfer | Llama + Transfer | |
Combined | 1.94 0.06 | 0.97 0.01 | 1.24 0.20 | 1.46 0.08 | 1.95 0.05 | 0.98 0.01 | 1.20 0.06 | 1.44 0.07 |
Burgers | 2.24 0.22 | 1.25 0.05 | 1.34 0.09 | 1.09 0.07 | 2.33 0.17 | 0.91 0.02 | 1.01 0.01 | 1.01 0.10 |
Heat | 1.01 0.16 | 0.56 0.04 | 0.63 0.00 | 0.83 0.03 | 0.91 0.04 | 0.43 0.05 | 0.52 0.02 | 0.67 0.24 |
Navier Stokes | 2.19 0.02 | 1.56 0.01 | 1.74 0.15 | 1.96 0.33 | 2.23 0.02 | 1.58 0.04 | 1.91 0.16 | 2.11 0.21 |
Shallow Water | 0.30 0.01 | 0.28 0.03 | 0.34 0.01 | 0.30 0.02 | 0.39 0.02 | 0.27 0.03 | 0.42 0.01 | 0.38 0.02 |
C.2 Additional Autoregressive Rollout Results
Table 4 below gives the numerical values for figure 9. These are the same results as in figure 4 with results for our combined data set included. Again, these results follow the trend of our next-step results closely.
Baseline | ST | Tokenizer | Llama | Transfer | ST + Transfer | Tokenizer + Transfer | Llama + Transfer | |
Combined | 0.24 0.02 | 0.06 0.00 | 0.14 0.06 | 0.23 0.02 | 0.24 0.02 | 0.06 0.00 | 0.11 0.03 | 0.23 0.02 |
Burgers | 0.19 0.18 | 0.05 0.01 | 0.09 0.06 | 0.05 0.00 | 0.33 0.26 | 0.04 0.00 | 0.06 0.02 | 0.04 0.01 |
Heat | 0.02 0.00 | 0.01 0.00 | 0.01 0.00 | 0.01 0.00 | 0.02 0.00 | 0.00 0.00 | 0.01 0.00 | 0.01 0.01 |
Navier Stokes | 0.58 0.02 | 0.13 0.00 | 0.22 0.10 | 0.37 0.22 | 0.59 0.04 | 0.13 0.01 | 0.28 0.09 | 0.48 0.23 |
Shallow Water | 49.44 4.73 | 54.28 15.15 | 45.15 12.36 | 49.42 7.35 | 108.57 43.56 | 37.32 4.40 | 47.20 31.23 | 30.70 10.47 |
C.3 Fixed-Future Additional Results
Table 5 below gives the numerical values for figure 10. These are the same results as in figure 5 with results for our combined data set included. We see the combined data set follows very similar trends to the Heat results, where Llama performs best but the other multimodal variants are close in performance.
Baseline | ST | Tokenizer | Llama | Transfer | ST + Transfer | Tokenizer + Transfer | Llama + Transfer | |
Combined | 0.527 0.004 | 0.334 0.009 | 0.362 0.058 | 0.283 0.015 | 0.527 0.004 | 0.332 0.008 | 0.361 0.058 | 0.270 0.009 |
Burgers | 0.486 0.003 | 0.431 0.004 | 0.341 0.007 | 0.317 0.015 | 0.486 0.003 | 0.409 0.009 | 0.309 0.016 | 0.247 0.018 |
Heat | 0.334 0.009 | 0.166 0.010 | 0.187 0.038 | 0.107 0.024 | 0.334 0.009 | 0.166 0.003 | 0.155 0.077 | 0.070 0.002 |
Navier Stokes | 0.674 0.008 | 0.353 0.020 | 0.390 0.042 | 0.352 0.011 | 0.674 0.008 | 0.379 0.021 | 0.447 0.083 | 0.383 0.029 |
Shallow Water | 0.004 0.001 | 0.003 0.000 | 0.004 0.000 | 0.003 0.000 | 0.004 0.001 | 0.004 0.000 | 0.004 0.001 | 0.003 0.000 |
C.4 Ablation Study Additional Results
Table 6 gives the numerical values for figure 11. The ablation study with Llama as the LLM backbone shows very similar to the SentenceTransformer backbone seen in figure 6. Adding qualitative information improves performance smaller than coefficient information, and only when coefficient information is not also present.
Equation | B | C | Q | BC | BQ | CQ | BCQ | |
Combined | 0.505 0.004 | 0.465 0.003 | 0.298 0.013 | 0.395 0.007 | 0.292 0.033 | 0.350 0.003 | 0.294 0.010 | 0.283 0.015 |
Burgers | 0.475 0.001 | 0.419 0.010 | 0.310 0.036 | 0.470 0.001 | 0.350 0.024 | 0.417 0.011 | 0.306 0.009 | 0.317 0.015 |
Heat | 0.319 0.007 | 0.244 0.006 | 0.159 0.013 | 0.221 0.013 | 0.162 0.050 | 0.186 0.030 | 0.155 0.012 | 0.107 0.024 |
Navier Stokes | 0.677 0.005 | 0.677 0.006 | 0.358 0.008 | 0.437 0.012 | 0.426 0.056 | 0.439 0.011 | 0.385 0.049 | 0.352 0.011 |
Shallow Water | 0.003 0.001 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.001 | 0.003 0.000 |
Table 7 gives the numerical values for figure 12. These are the same results as in figure 6 with results from our combined data set added. Again, we see a small improvement in performance when qualitative information is added in tandem with coefficient information in our Burgers and combined data sets, however this is not present in our Navier Stokes data set.
Equation | B | C | Q | BC | BQ | CQ | BCQ | |
Combined | 0.502 0.004 | 0.474 0.003 | 0.347 0.002 | 0.396 0.005 | 0.364 0.012 | 0.364 0.008 | 0.326 0.005 | 0.332 0.008 |
Burgers | 0.475 0.001 | 0.435 0.003 | 0.387 0.003 | 0.471 0.001 | 0.412 0.005 | 0.434 0.010 | 0.378 0.001 | 0.429 0.002 |
Heat | 0.319 0.008 | 0.268 0.004 | 0.251 0.005 | 0.221 0.014 | 0.256 0.015 | 0.171 0.007 | 0.198 0.012 | 0.165 0.011 |
Navier Stokes | 0.676 0.003 | 0.675 0.003 | 0.352 0.012 | 0.460 0.039 | 0.352 0.016 | 0.444 0.021 | 0.353 0.015 | 0.353 0.020 |
Shallow Water | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 | 0.003 0.000 |
Appendix D Rollout
@@
Figure 13 plots the autoregressive rollout error against timestep. Errors are first averaged over each sample for a given random seed, then the mean and standard deviation across seeds is plotted.
Appendix E Additional t-SNE Results
@@
We see as we increase sentence complexity, we get additional structure in our t-SNE embeddings. Adding coefficient information captures the distribution of coefficients well, adding boundary conditions captures the distinct categories well. Using both coefficient information and boundary condition information shows clustering by boundary condition, as well as more clusters than when using qualitative information. These results are true for both Llama embeddings in figures 14 and 15, as well as SentenceTransformer embeddings in figures 16 and 17.